# Noé Arnold, Juan Beccuti, Thunj Chantramonklasri, Matthias Hafner, Nicolas Oderbolz Center for Cryptoeconomics

### v1

@@ III-A Baseline Model
In addition, ETH holders who stake through an intermediary incur fees, which are modeled in the following as a tax on staking revenues and as such are treated separately from the cost functions. The fee rate for the cSSP is denoted as $f_{c}$ and the fee rate for the dSSP is denoted as $f_{d}$, both of which are assumed to be exogenous in the baseline model.
Under the assumption of a constant ETH price, thus isolating staking decisions from price volatility and inflation effects, each staker maximizes expected annual profit. Experts choose a deposit $\hat{d}_{ss}\geq 0$ to maximize
$\max_{\hat{d}_{ss}}\,\Big{\{}y_{i}(D)\cdot\hat{d}_{ss}+y_{v}\cdot P_{ss}(D)-\Big{(}C_{ss}+c_{ss}\cdot\hat{d}_{ss}^{\alpha_{ss}}\Big{)}\Big{\}},$ (1)
where
@@
and $P_{ss}(D)$ is the probability that a Solo Staker choosing $\hat{d}_{ss}$ becomes a proposer
$P_{ss}(D)=\frac{\hat{d}_{ss}}{\hat{d}_{ss}+(N_{ss}-1)\cdot d_{ss}+N_{r}\cdot d_{r}+N_{t}\cdot d_{t}}$
@@
Techies choose a deposit $\hat{d}_{t}\geq 0$ to maximize
$\max_{\hat{d}_{t}}\,\Big{\{}(1-f_{d})\cdot\Big{[}y_{i}(D)\cdot\hat{d}_{t}+y_{v}\cdot P_{t}(D)\Big{]}+y_{d}\cdot\hat{d}_{t}$
$-\Big{(}C_{t}+c_{t}\cdot\hat{d}_{t}^{\alpha_{t}}\Big{)}\Big{\}},$ (2)
with
$P_{t}(D)=\frac{\hat{d}_{t}}{N_{ss}\cdot d_{ss}+\hat{d}_{t}+(N_{t}-1)\cdot d_{t}+N_{r}\cdot d_{r}}.$
Homogeneity again implies $\hat{d}_{t}^{*}=d_{t}$ in equilibrium.
Finally, Retailers choose a deposit $\hat{d}_{r}\geq 0$ to maximize
$\max_{\hat{d}_{r}}\,\Big{\{}(1-f_{c})\cdot\Big{[}y_{i}(D)\cdot\hat{d}_{r}+y_{v}\cdot P_{r}(D)\Big{]}$
$-\Big{(}C_{r}+c_{r}\cdot\hat{d}_{r}^{\alpha_{r}}\Big{)}\Big{\}},$ (3)
with
$P_{r}(D)=\frac{\hat{d}_{r}}{N_{ss}\cdot d_{ss}+N_{t}\cdot d_{t}+\hat{d}_{r}+(N_{r}-1)\cdot d_{r}}.$
ETH holders act strategically, anticipating the staking decisions of other market participants when determining their own profit-maximizing level of stake. The resulting model solution thus represents a Nash-equilibrium. We refer to Section -A of the Appendix for a detailed derivation.
III-B Model Extensions
We extend the baseline game theoretic model by sequentially introducing important refinements to the characteristics of stakers. First, we introduce solo stakers who face a larger variance in MEV payoffs than other stakers. Next, we relax the assumption that all stakers are fully strategic by introducing inattentive agents who do not respond to changes in the issuance curve. Finally, we introduce an intermediary dSSP with market power.
III-B1 Variability in MEV rewards for solo stakers
In the baseline model, expected MEV revenues are assumed to be identical whether an ETH holder chooses solo staking or joins a pool through an SSP. However, since MEV revenues can vary substantially over time, this formulation does not account for the fact that solo stakers may take several years to propose a block with high MEV, unlike stakers in a pool who can benefit from smoothing MEV revenues over time. To account for such a difference, we introduce the variance over MEV revenues for a (moderately risk-averse) solo staker, while assuming that this variance can be neglected for the other types, i.e., that pools are large enough. Thus, the maximization problem of a solo staker (assumed to be of type Expert) becomes [^1] See Section -B of the Appendix for the derivation.
$\max_{\hat{d}_{e}}\quad y_{i}(D)\cdot\hat{d}_{e}+[y_{v}\cdot P_{ss}(D)-y_{v}^{2}\cdot P_{ss}(D)^{2}]$
$-(C_{ss}+c_{ss}\cdot\hat{d}_{e}^{\alpha_{e}}),$ (4)
@@ III-B1 Variability in MEV rewards for solo stakers
III-B2 Inattentive Agents
Further, we divide the Retailer category into two subgroups: (1) institutions and (2) inattentive retail investors. The latter group is assumed to be unresponsive to changes in the issuance yield. This behavior could be interpreted as the agents being inattentive to changes in issuance or that their profit-maximizing decisions result in a corner solution, as is the case in the theoretical results obtained by Ma, (2024). The primary purpose of incorporating this type of agent into the model is to assess how their presence influences the staking behavior of other types.
III-B3 Intermediary dSSP with Market Power
@@
We extend the model to capture such interactions between an intermediary, stakers, and multiple centralized staking service providers (cSSPs). In doing so, our framework specifically captures Lido, the dominant player in this intermediation space. Instead of assuming perfect competition, we model the decentralized staking service provider (dSSP) as a player with market power. We incorporate this extension in addition to the previously described extensions that account for inattentive retail stakers and the higher variance in execution rewards faced by solo stakers.
The goal is to examine how changes in the issuance curve affect the dSSP’s fee-setting decisions and, consequently, overall staking levels. We maintain the timing structure from previous model iterations: the dSSP sets its fees first, followed by the stakers setting their deposit levels.
@@
We maintain the assumption of a segregated staking market and assume that the ETH holder type Techie has access to choose the dSSP as an intermediary. In doing so, they are indifferent to the specifics of the dSSP-cSSP relationship; their decision is based solely on Lido’s fees ($f_{d}$), their expected rewards, and associated costs. Essentially, they solve the same optimization problem as in equation (2): each Techie selects $\hat{d}_{t}\geq 0$ to maximize their profit
$\max_{\hat{d}_{t}}\Big{\{}(1-f_{d})\cdot\left[y_{i}(D)\cdot\hat{d}_{t}+y_{v}\cdot P_{t}(D)\right]$
$+y_{d}\cdot\hat{d}_{t}-(C_{t}+c_{t}\cdot\hat{d}_{t}^{\alpha_{t}})\Big{\}},$ (5)
To maximize its profits, the dSSP (Lido) sets both the fee $f_{d}$ charged to Techies and the fee $\hat{f}_{d}$ paid to cSSPs. The dSSP holds the bargaining power to determine these fees, as there are many cSSPs competing to be integrated into the platform. We assume that the dSSP charges a uniform fee to each cSSP and distributes stakes evenly among them. The profit maximization problem of the dSSP is therefore:
$\max_{f_{d},\hat{f}_{d}}$$\Big{\{}(f_{d}-\hat{f}_{d})\cdot$
$\left[y(D_{-t},D_{t}(f_{d}))\cdot D_{t}(f_{d})+y_{v}(D_{-t},D_{t}(f_{d}))\right]$
$-C_{i}(D_{t}(f_{d}))\Big{\}},$ (6)
@@
where $D_{t}=\sum\hat{d}_{t}$, and $D_{-t}$ denotes the stakes from non-Techies. We assume that Lido enjoys economies of scale, i.e., $C_{i}(D_{t})$ is a concave function. In setting $\hat{f}_{d}$, Lido ensures that cSSP profits, $\pi_{cSSP}$, remain at or above zero. In other words, Lido considers the cSSPs’ participation constraints (PC) to secure sufficient cSSP participation for allocating Techies’ stakes.
Consider any cSSP $m=1...M$. This cSSP sets the fee $f_{c}^{m}$ charged to stakers who bypass the dSSP (Lido) and stake directly with it (see Staker 1 in Figure 1).
$\max_{f_{c}^{m}}\quad f_{c}^{m}\cdot\left[y(D)\cdot D_{i}^{m}+y_{v}(D_{i}^{m},D)\right]$
$+\hat{f}_{d}\cdot\left[y(D)\cdot D_{t}^{m}+y_{v}(D_{t}^{m},D)\right]$
$-C_{c}(D_{t}^{m}+D_{i}^{m}),$ (7)
@@ IV Results
where $D$ represents the staking deposit level. The second issuance schedule follows one of the alternative proposals presented by Elowsson, (2024):
@@
IV-A Model Simulation
As a first step, the general relationships between the model parameters and the resulting equilibrium outcomes are explored. The goal is to develop an intuition for the relevant factors driving equilibrium staking decisions and how they might drive differences in the outcomes under the two consensus issuance schedules. This is done by running simulations of the model, randomly varying the parameter values in each simulation run. We conduct 5 million simulation runs for each of the two issuance schedules.
@@ IV-A Model Simulation
  • Observation 1: Higher marginal costs lead to smaller adjustments in the equilibrium staking supply when transitioning from $y_{i}(D)$ to $y^{\prime}_{i}(D)$. This effect is illustrated in Figures 3(a) - 3(c), using solo staking as an example.
  • Observation 2: When stakers have access to additional MEV revenues or DeFi yields, they exhibit a weaker response in their equilibrium staking supply when transitioning from $y_{i}(D)$ to $y^{\prime}_{i}(D)$. Figures 3(d) and 3(e) illustrate this effect using staking via dSSP as an example.
@@
(a) $D_{ss}$ vs. $C_{ss}$
(b) $D_{ss}$ vs. $c_{ss}$
(c) $D_{ss}$ vs. $\alpha_{ss}$
(d) $D_{dSSP}$ vs. $y_{v}$
(e) $D_{dSSP}$ vs. $y_{d}$
(f) $D_{ss}$ vs. $y_{d}$
@@
Recall that the empirical evidence discussed in Section II suggests that solo stakers tend to be more sensitive to changes in staking yield than other staking categories. Within the framework of the model, an increased sensitivity of solo stakers could arise from several factors. One possible explanation is that the marginal cost of solo staking is lower relative to other staking methods, making solo stakers more responsive to fluctuations in rewards. Alternatively, solo stakers may have more limited access to external revenue opportunities or may discount expected MEV revenues more than other market participants due to their variability. Finally, since ETH holders compete for consensus issuance yields and we assume strategic solo stakers, either of these dynamics could be exacerbated by other types of ETH holders being relatively increasingly less elastic to changes in consensus issue yields.
IV-B Comparative Analysis
Below we analyze the equilibrium results of the game theoretic model based on specific parameter calibrations. The calibration of the model parameters, in particular the underlying assumptions about cost structures and revenue streams, is described in detail in Section -C of the Appendix. In essence, the calibration approach relies on setting cost parameters such that, under the current issuance schedule, the model produces today’s distribution of stake offerings. However, the calibrated model may not reflect the true population parameters, especially if the model does not capture all of the factors that drive agents’ decisions in the staking market.
IV-B1 Baseline Model

### v2

@@ 4.4 Staking Costs
Later, we calibrate the model looking for the set of parameters that gives a similar distribution of staking supply than the one observed in the market at the moment of this study. We also assume that the costs functions are such that there are interior solutions. [^6] This assumption is standard in models where agents face no binding constraints on their choices, such as budget constraints. The assumptions made when calibrating the cost parameters are described in Section 4.7.2.
@@
Furthermore, ETH holders staking through an intermediary incur fees, which are modeled as a tax on revenues and as such treated separately from the cost functions. The cSSP fee rate is denoted as $f_{c}$, and the dSSP fee rate as $f_{d}$, both of which are assumed to be exogenous in the baseline model. As such, fees are simply assumed to be fixed and not determined by the strategic behavior of the staking service providers in the market. Subsequent extensions will relax this assumption and to some extent allow fees to be determined as a function of the profit maximization problem of the staking service providers in the model. Again, the assumptions made when calibrating the fee parameters are described in Section 4.7.2.
4.5 Agent Staking Decisions
Each ETH holder strategically decides how much to stake in order to maximize their profits, taking into account the staking behavior of others. That is, stakers anticipate the actions of other participants and their effects on the inflationary reward, and adjust their own staking levels accordingly to achieve a profit-maximizing outcome. The resulting model solution thus represents a Nash-equilibrium. We refer to Section A of the Appendix for a detailed derivation.
The following model can be interpreted as one in which a staker is endowed with a fixed amount of fiat currency and must decide how much to allocate to ETH staking versus how much to consume or use for other purposes. To simplify the analysis, we introduce several assumptions. The opportunity cost of foregone consumption is implicitly captured by a convex cost function. [^7] This also justifies the use of a convex cost function, as it reflects the increasing marginal cost of substituting away from alternative consumption or uses. As a result, the staker’s utility (or profit) function is concave, leading to interior solutions to the maximization problem—even in the absence of binding budget constraints, as we assume here.
For simplicity, we abstract from fluctuations in the ETH price, including expectations about future price movements and any associated capital gains or losses. That is, we assume a constant ETH price, thereby isolating the staking decision from market volatility and the negative effects of inflation on the token price. In addition, stakers are assumed to not discount the future. Under all these assumptions, the staker’s problem reduces to maximizing expected annual profit. We next describe each staker type decision problem:
Given the deposits of other stakers $\{d_{ss},d_{r},d_{t}\}$, a representative Retailer chooses a deposit $\hat{d}_{r}\geq 0$ to maximize
$\Pi_{r}\equiv$ $\max_{\hat{d}_{r}}\,\Big{\{}(1-f_{c})\cdot\Big{[}y_{i}(D)\cdot\hat{d}_{r}+y_{v}\cdot P_{r}(D)\Big{]}-\Big{(}C_{r}+c_{r}\cdot\hat{d}_{r}^{\alpha_{r}}\Big{)}\Big{\}},$ (1)
such that:
$\Pi_{r}\geq 0$
$D=N_{ss}\cdot d_{ss}+N_{t}\cdot d_{t}+\hat{d}_{r}+(N_{r}-1)\cdot d_{r}\in[0,120M],$
$P_{r}(D)=\frac{\hat{d}_{r}}{N_{ss}\cdot d_{ss}+N_{t}\cdot d_{t}+\hat{d}_{r}+(N_{r}-1)\cdot d_{r}}\in[0,1],$
where $f_{c}$ is the fee charged by $cSSPs$.
Let $\hat{d}_{r}^{*}$ denote the optimal choice of the maximization problem. Due to homogeneity within each staker type category, all Retailers must choose the same deposit level in equilibrium. That is, in equilibrium, the individual optimal choice $\hat{d}_{r}^{*}$ satisfies $\hat{d}_{r}^{*}=d_{r}$.
$P_{r}(D)$ denotes the probability that a staker with deposit $\hat{d}_{r}$ becomes a proposer. We model this probability as the ratio of the staker’s deposits to the total deposits, as indicated in the denominator. A more accurate representation would model this as the ratio of the number of validators controlled by the staker to the total number of validators. This would require a step function that increments with every 32 ETH, corresponding to the validator activation threshold. For tractability, we instead model $P_{r}$ as a continuous function of $\hat{d}_{r}$, which simplifies the analysis.
Similarly, each Techie chooses a deposit $\hat{d}_{t}\geq 0$ to maximize
$\Pi_{t}\equiv$ $\max_{\hat{d}_{t}}\,\Big{\{}(1-f_{d})\cdot\Big{[}y_{i}(D)\cdot\hat{d}_{t}+y_{v}\cdot P_{t}(D)\Big{]}+y_{d}\cdot\hat{d}_{t}-\Big{(}C_{t}+c_{t}\cdot\hat{d}_{t}^{\alpha_{t}}\Big{)}\Big{\}},$ (2)
such that:
$\Pi_{t}\geq\Pi_{r},$
$D=N_{ss}\cdot d_{ss}+\hat{d}_{t}+(N_{t}-1)\cdot d_{t}+N_{r}\cdot d_{r}\in[0,120M],$
$P_{t}(D)=\frac{\hat{d}_{t}}{N_{ss}\cdot d_{ss}+\hat{d}_{t}+(N_{t}-1)\cdot d_{t}+N_{r}\cdot d_{r}}\in[0,1].$
Homogeneity again implies $\hat{d}_{t}^{*}=d_{t}$. Notice that this type of staker pays a fee $f_{d}$ and enjoys of DeFi yields $y_{d}$.
A representative Expert chooses a deposit $\hat{d}_{ss}\geq 0$ to maximize
$\Pi_{ss}\equiv$ $\max_{\hat{d}_{ss}}\,\Big{\{}y_{i}(D)\cdot\hat{d}_{ss}+y_{v}\cdot P_{ss}(D)-\Big{(}C_{ss}+c_{ss}\cdot\hat{d}_{ss}^{\alpha_{ss}}\Big{)}\Big{\}},$ (3)
such that:
$\Pi_{ss}\geq 0$
@@ 4.5 Agent Staking Decisions
$P_{ss}(D)=\frac{\hat{d}_{ss}}{\hat{d}_{ss}+(N_{ss}-1)\cdot d_{ss}+N_{r}\cdot d_{r}+N_{t}\cdot d_{t}}\in[0,1].$
@@
4.6 Model Extensions
We extend the baseline game-theoretic model by incorporating key refinements to the characteristics of staking agents, building each extension sequentially. First, we introduce solo stakers facing greater variance in MEV revenues than other stakers. Next, we relax the assumption that all stakers are fully strategic by introducing inattentive agents who do not respond to changes in the issuance curve. Finally, we enhance the model of the intermediary market by incorporating a dSSP with market power.
4.6.1 Variability in MEV rewards for solo stakers
In the baseline model, expected MEV revenues are assumed to be identical regardless of whether an ETH holder chooses to solo stake or joins a pool via a staking service provider. However, since MEV revenues can vary substantially over time, this formulation does not account for the fact that solo stakers may require several years to propose a block with high MEV, unlike other stakers in a pool who can benefit from smoothing MEV revenues over time. To incorporate such difference, we introduce the variance over such reward (denoted below as $Y_{MEV}$) for a moderately risk-averse solo stakers, while we assume that for the other types that variance can be neglected, i.e., pools are large enough. Hence, the maximization problem of a solo staker (assumed to be of type Expert) becomes: [^8] See the Appendix for the derivation and discussion.
$\max_{\hat{d}_{e}}\quad y_{i}(D)\cdot\hat{d}_{e}+[y_{v}\cdot P_{ss}(D)-y_{v}^{2}\cdot P_{ss}(D)^{2}]-(C_{ss}+c_{ss}\cdot\hat{d}_{e}^{\alpha_{e}}),$ (4)
@@ 4.6.1 Variability in MEV rewards for solo stakers
4.6.2 Inattentive Agents
Further, we divide the Retailer category into two subgroups: (1) institutions and (2) inattentive retail investors. The latter group is assumed to be unresponsive to changes in the issuance yield. This behavior could be interpreted as the agents being inattentive to changes in issuance or that their profit-maximizing decisions result in a corner solution, as is the case in the theoretical results obtained by Ma, (2024). [^9] Evidence from other blockchains suggests that the presence of inattentive stakers may be a relevant phenomenon. For instance, Liu et al., (2024) find that approximately $80\%$ of delegator accounts on the Cardano network are ”sleepy” delegators, representing around $50\%$ of the total staked volume. An interesting research direction is to study how prevalent inattentive staking behavior is within the Ethereum ecosystem. For our analysis, we assume that only $20\%$ of the volume staked belongs to inattentive stakers. The primary purpose of incorporating this type of agent into the model is to assess how their presence influences the staking behavior of other types.
4.6.3 Intermediary dSSP with Market Power
@@
In the following, we extend our model to incorporate the interactions between an intermediary, stakers, and multiple centralized staking service providers (cSSPs). Our framework specifically captures Lido, the dominant player in this intermediation space. Rather than assuming perfect competition, we model the decentralized staking service provider (dSSP) as a player with market power, building on previous extensions that account for inattentive retail stakers and the higher variance in execution rewards faced by solo stakers.
Our objective is to examine how changes in the issuance curve influence the dSSP’s fee-setting decisions and, consequently, the overall staking level. We preserve the timing structure from earlier model iterations: the dSSP sets its fees first, followed by stakers determining their deposit levels.
@@
  • Techies: We maintain the assumption of a segmented staking market and assume that the ETH holder type Techie chooses the dSSP as an intermediary. In doing so, they are indifferent to the specifics of the dSSP-cSSP relationship; their decision is based solely on Lido’s fees ($f_{d}$), their expected rewards, and associated costs. Essentially, they solve the same optimization problem as in equation (2): each Techie selects $\hat{d}_{t}\geq 0$ to maximize their profit
    $\max_{\hat{d}_{t}}\quad(1-f_{d})\cdot\left[y_{i}(D)\cdot\hat{d}_{t}+y_{v}\cdot P_{t}(D)\right]+y_{d}\cdot\hat{d}_{t}-(C_{t}+c_{t}\cdot\hat{d}_{t}^{\alpha_{t}}),$ (5)
  • Intermediary (Lido): To maximize its profits, the dSSP sets both the fee $f_{d}$ charged to Techies and the fee $\hat{f}_{d}$ paid to cSSPs. The dSSP holds the bargaining power to determine these fees, as there are many cSSPs competing to be integrated into the platform. We assume that the dSSP charges a uniform fee to each cSSP and distributes stakes evenly among them. The profit maximization problem of the dSSP is therefore:
    $\max_{f_{d},\hat{f}_{d}}$ $\quad(f_{d}-\hat{f}_{d})\cdot\left[y(D_{-t},D_{t}(f_{d}))\cdot D_{t}(f_{d})+y_{v}(D_{-t},D_{t}(f_{d}))\right]-C_{I}(D_{t}(f_{d})),$ (6)
@@
    where $D_{t}=\sum\hat{d}_{t}$ denotes the total stakes by Techies, and $D_{-t}$ the stakes from non-Techies. We assume that the dSSP enjoys economies of scale, i.e., $C_{I}(D_{t})$ is a concave function. In setting $\hat{f}_{d}$, the dSSP ensures that cSSP profits, $\pi_{cSSP}$, remain at or above zero. In other words, the dSSP considers the cSSPs’ participation constraints (PC) to secure sufficient cSSP participation for allocating Techies’ stakes.
  • cSSP: Consider any cSSP $m=1...M$. This cSSP sets the fee $f_{c}^{m}$ charged to stakers who bypass the dSSP and stake directly with it (e.g., see Staker 1 in Figure 4).
    $\max_{f_{c}^{m}}$ $\quad f_{c}^{m}\cdot\left[y(D)\cdot D_{i}^{m}+y_{v}(D_{i}^{m},D)\right]+\hat{f}_{d}\cdot\left[y(D)\cdot D_{t}^{m}+y_{v}(D_{t}^{m},D)\right]$ (7)
[…]
@@ 4.7 Results
where $D$ represents the staking deposit level. The second issuance schedule follows one of the alternative proposals presented by Elowsson, 2024c :
@@
4.7.1 Numerical Simulation
As a first step, the general relationships between the model parameters and the resulting equilibrium outcomes are explored. The goal is to develop an intuition for the relevant factors driving equilibrium staking decisions and how they drive differences in the outcomes under the two consensus issuance schedules. This is done by running simulations of the model, randomly varying the parameter values in each simulation run. [^10] To account for the large parameter space, we conduct 5 million simulation runs for each of the two issuance schedules.
@@ 4.7.1 Numerical Simulation
  • Observation 1: Everything else held equal, a staking solution having higher marginal costs is associated with smaller adjustments in the equilibrium staking supply when transitioning from $y_{i}(D)$ to $y^{\prime}_{i}(D)$. This effect is illustrated in Figures 5(a) - 5(c), using solo staking as an example.
  • Observation 2: Stakers having access to additional MEV revenues or DeFi yields is associated with a weaker response in their equilibrium staking supply when transitioning from $y_{i}(D)$ to $y^{\prime}_{i}(D)$. Figures 5(d) and 5(e) illustrate this effect using staking via dSSP as an example.
@@
((a)) $D_{ss}$ vs. $C_{ss}$
((b)) $D_{ss}$ vs. $c_{ss}$
((c)) $D_{ss}$ vs. $\alpha_{ss}$
((d)) $D_{dSSP}$ vs. $y_{v}$
((e)) $D_{dSSP}$ vs. $y_{d}$
((f)) $D_{ss}$ vs. $y_{d}$
@@
4.7.2 Model Calibration
We calibrate the baseline model with specific parameters settings and compare Nash equilibria under two issuance schedules. Given the number of parameters to be calibrated, we find that estimating model parameters based on a method of simulated moments, i.e., algorithmically setting parameters to minimize the distance between simulated and observed market shares, does not provide a stable best fit. In other words, we find that a variety of possible parameter combinations can yield simulated market shares that are consistent with the observed data. This finding is illustrated in Figure 6. For different combinations of parameter settings for $c_{t}$ and $c_{r}$, we find that a range of parameter settings are able to minimize the distance between simulated market shares and observed market shares. This issue is exacerbated when allowing the algorithm to search freely over an increasing number of parameters.
Figure 6: Median distance (L2 norm) of simulated market shares compared to observed market shares for different values of $c_{t}$ and $c_{r}$
Note: This plot displays the median distance (L2 norm) between simulated market shares and observed market shares for different values of $c_{t}$ and $c_{r}$.
As an alternative approach, we propose parameter settings that are broadly consistent with the available literature and our qualitative understanding of the Ethereum staking market and that in the baseline model, achieve a close fit to the observed market shares (see Table 1). Additionally, we conducted a survey on the Reddit forum ethstaker, focusing on the cost structure and decisions faced by solo stakers. [^11] See https://www.reddit.com/r/ethstaker/comments/1dneiy3/solo_stakers_questionnaire/
Table 1: Reference staking distribution used in parameter calibration
Staking Method M ETH %
dSSP $18$ $54.1\%$
cSSP $14.4$ $44.6\%$
Solo Staking $0.9$ $2.7\%$
TOTAL $34.2$
Note: The estimated number of validators is based on a simplification of the categorization reported by (Kotelskiy et al.,, 2024) and is calculated as the amount of staked ETH divided by 32. We categorize DSMs, LRTs, Whales, and Unidentified as dSSPs. Meanwhile, cSSPs consist of CSPs and CeXs.
First, we assume that solo staking incurs fixed costs, while staking via a dSSP or cSSP does not:
$C_{ss}=0.4,\quad C_{t}=0,\quad C_{r}=0.$
Contrary to staking through an intermediary, solo staking requires an individual to set up an Ethereum validator node. To do this, solo stakers can purchase dedicated hardware, subscribe to a cloud computing service, or use existing hardware. The costs associated with each of these options vary substantially. To further complicate matters, a range of hardware specifications can be used to run a node, each with varying costs. To date, the literature has made highly simplifying assumptions about the hardware costs of solo validators. In particular, both pa7x1, (2024) and Kotelskiy et al., (2024) assume that a solo staker factors in 1000 USD to initially purchase the necessary hardware, which they amortize over a period of 5 years. The anecdotal evidence collected from the survey on ethstaker generally supports this assumption but also reveals substantial variation owing to the fact that some stakers use existing hardware and do not purchase dedicated hardware for staking, i.e. likely treat their hardware expenditure as sunk costs.
Similar to hardware costs, we assume costs for energy and internet to be fixed, since they are a fundamental requirement to running an Ethereum node but do not scale when adding additional validators to a node. Again, pa7x1, (2024) and Kotelskiy et al., (2024) assume around 1.4 to 2 USD in energy costs per week and around 0 to 12 USD per week for internet costs. These assumptions are generally supported by the anecdotal evidence from the ethstaker survey.
Taken together, we assume fixed costs of $\approx 1^{\prime}000\$/year$. This includes hardware which is usually assumed to be amortized in $5$ years ($200-400$$\$/year$), high-speed and stable internet connection ($50$$\$/month$), and additional electricity expenditures ($100$$\$/year$). Denoting this cost in terms of ETH, at a price of ETH in USD of 2’400 USD, we obtain a parameter setting of $C_{ss}=0.4$. Again, we assume that staking through an intermediary does not create any fixed costs from the perspective of the ETH holder. Therefore, we set $C_{r}=0$ and $C_{r}=0$.
Further, we are argue the following general assumptions to hold regarding variable costs
$c_{ss}<c_{t}<c_{r}\quad\text{and}\quad\alpha_{ss}>\alpha_{t}=\alpha_{r}.$
The literature is quite sparse on variable staking costs. As a result, we rely primarily on anecdotal evidence to support these assumptions. We argue that imposing the above conditions may be reasonable when considering the maintenance costs of solo staking. As the number of validators increases, complexity and operational overhead may increase non-linearly. For example, more validators require better network management. Managing and updating multiple nodes, as well as handling security patches and upgrades, requires more effort as the infrastructure scales. Based on the results of the ethstaker survey, it is also plausible to assume that, at least at lower levels of staking, stakers are hobbyists and do not price in maintenance and upkeep costs. As staking operations increase in size, we expect that individual stakers begin to price in maintenance and upkeep costs when choosing their optimal.
Moreover, we assume that the variable costs for staking via a cSSP exceed those for staking via a dSSP. This assumption is based on the observation that, in addition to the inherent risks of staking (e.g., slashing risks), centralized exchanges introduce additional custodial risks—risks that are absent in decentralized staking service providers such as Lido, which do not take custody of users’ assets. In addition, institutional stakers staking via a centralized solution may accrue additional legal costs. Finally, we assume that $\alpha_{j}>1$ for all staking methods $j$. This allows for convex cost functions which capture the fact that ETH holders may be resource constrained. This assumption is standard in models where agents face no binding constraints on their choices, such as budget constraints.
Overall, due to the argumentative nature of our calibration approach, we argue that the results presented below should not be interpreted as predictive point estimates. Rather, the focus should be on the relative changes and dynamics observed between the Nash equilibria under the two issuance schedules being considered.
Table 2: Summary of parameter settings
Parameter Description Parameter Value
$N_{ss}$ Number of agents of type Expert $N_{ss}=25^{\prime}000$
$N_{t}$ Number of agents of type Techie $N_{t}=200^{\prime}000$
$N_{r}$ Number of agents of type Retailer $N_{r}=925^{\prime}000$
$y_{v}$ Annual MEV revenue $y_{v}=300^{\prime}000$
$y_{d}$ Annual DeFi yield $y_{d}=0.02$
$f_{d}$ Fee dSSP $f_{d}=0.1$
$f_{c}$ Fee cSSP $f_{c}=0.25$
$C_{ss}$ Fixed cost parameter solo staking $C_{ss}=0.4$
$c_{ss}$ Variable cost parameter solo staking $c_{ss}=0.00053$
$\alpha_{ss}$ Exponent cost function solo staking $\alpha_{ss}=2$
$C_{t}$ Fixed cost parameter for Techie staking via dSSP $C_{t}=0$
$c_{t}$ Variable cost parameter for Techie staking via dSSP $c_{t}=0.0038$
$\alpha_{t}$ Exponent cost function for Techie staking via dSSP $\alpha_{t}=1.5$
$C_{r}$ Fixed cost parameter for Retailer staking via cSSP $C_{r}=0$
$c_{r}$ Variable cost parameter for Retailer staking via cSSP $c_{r}=0.0048$
$\alpha_{r}$ Exponent cost function for Retailer staking via cSSP $\alpha_{r}=1.5$
4.7.3 Comparative Analysis
Below we analyze the equilibrium results of the game-theoretic model based on specific parameter calibrations described in the previous section. Effectively, we compare the equilibrium staking outcomes of a calibrated baseline model with the equilibrium outcomes of a calibrated model under the alternative issuance schedule $y^{\prime}_{i}(D)$.
Baseline Model